As you may have seen - the data from last night exhibits problems. This was due to mistake in tuning the timing (by me) for the inel counters early last evening. Despite the change looked reasonable (with scope, rates etc) in was most likely distorted by a very bad fill /steering. This morning DB and I checked with a pulser source the timing and basically reverting to the settings before yesterday is the correct thing to do. Thus unfortunately the data taken from 1/8-1/9 should be discarded i.e. runs 6347 - 6388 and the program should be repeated. (Completion of 12 deg 1/4A pol etc..) Also for the runs 6322-6346 in particular the FS data are of little value due to the HV problems that caused frequent trips of the front trigger slats. (solved yesterdday morning). PS Glad the shift crew did not call about the web-shift page - I have no clues on solving this problem. ------------------------------------------------------ Flemming Videbaek Physics Department Brookhaven National Laboratory tlf: 631-344-4106 fax 631-344-1334 e-mail: videbaek@bnl.gov ----- Original Message ----- From: "Apache" <apache@pii3.brahms.bnl.gov> To: <brahms-l@bnl.gov>; <olchansk@bnl.gov>; <videbaek@bnl.gov> Sent: Tuesday, January 08, 2002 11:30 PM Subject: Shift report 20020108 16:00-24:00 > > ------------------------------------------------------------ > Date: Tue Jan 8 23:30:22 2002, Shift: 16:00-24:00 > Supervisor: IGB, with: mads > Log book#12 pages 29 - 30 > > Shift Summary: > ------------------------------------------------------------ > Machine Status: > Had one store from 1730 until 2130. The beam quality was not > terribly good, but as it was impossible to get a new fill, > decided that not good beam was still better than no beam. > No beam since 2130, but the machine is being filled for a ramp. > > Achievements: > Finished 100k events with the inner innelastic counters out > of the trigger. > Replaced these signals in the discriminator > > Detector Problems: > > Flemming fixed a small timing problem in the trigger, > otherwise no problems were noticed. > > We were unable to submit the shift check using the Web form. > We decided that it was not worth interupting Flemmings > (late, because of the timing problem) supper for > this minor problem. > > Had some problems getting the DAQ to work properly. > FV fixed this, though I'm not completely sure how, exactly, he managed > this. > ------------------------------------------------------------ > ------------------------------------------------------ Flemming Videbaek Physics Department Brookhaven National Laboratory tlf: 631-344-4106 fax 631-344-1334 e-mail: videbaek@bnl.gov
This archive was generated by hypermail 2b30 : Wed Jan 09 2002 - 11:09:20 EST